OpenStudy (anonymous):

what is the value of the coefficiant b in the quadratic equation y=(2x+1) (4x-3) I got -2 am I correct

OpenStudy (jdoe0001):

2x * (-3) = -6x +1 * (4x) = 4x -6x + 4x = -2x so, yes, the coefficient is "-2"